# Diffusion-based Task-oriented Semantic Communications with Model Inversion Attack

### v1

@@ Abstract
Semantic communication has emerged as a promising neural network-based system design for 6G networks. Task-oriented semantic communication is a novel paradigm whose core goal is to efficiently complete specific tasks by transmitting semantic information, optimizing communication efficiency and task performance. The key challenge lies in preserving privacy while maintaining task accuracy, as this scenario is susceptible to model inversion attacks. In such attacks, adversaries can restore or even reconstruct input data by analyzing and processing model outputs, owing to the neural network-based nature of the systems. In addition, traditional systems use image quality indicators (such as PSNR or SSIM) to assess attack severity, which may be inadequate for task-oriented semantic communication, since visual differences do not necessarily ensure semantic divergence. In this paper, we propose a diffusion-based semantic communication framework, named DiffSem, that optimizes semantic information reconstruction through a diffusion mechanism with self-referential label embedding to significantly improve task performance. Our model also compensates channel noise and adopt semantic information distortion to ensure the robustness of the system in various signal-to-noise ratio environments. To evaluate the attacker’s effectiveness, we propose a new metric that better quantifies the semantic fidelity of estimations from the adversary. Experimental results based on this criterion show that on the MNIST dataset, DiffSem improves the classification accuracy by $10.03\%$, and maintain stable performance under dynamic channels. Our results further demonstrate that significant deviation exists between traditional image quality indicators and the leakage of task-relevant semantic information.
Index Terms:
Semantic communications, diffusion model, privacy preserving, model inversion attack.
@@ I Introduction
Deep learning based semantic communication system has emerged as a promising solution to enhance information transmission efficiency in sixth-generation (6G) wireless networks. Unlike traditional communication paradigms that prioritize the symbol-level accurate transmission, semantic communication shifts the focus to conveying the intended meaning through symbols [1]. In such systems, the transmitter and receiver collaboratively design mappings between meanings and channel symbols, naturally aligning with the principles of joint source-channel coding (JSCC) [2]. This approach accommodates more complex data sources (e.g., images) and diverse distortion measures, such as classification loss [3] or task-oriented metrics [4]. As a result, semantic communication has been studied for various types of sources such as text [5, 6], visual data [7, 8, 9] and video [10, 11], and has also been applied in edge computing [12, 13] and autonomous driving [14, 15]. By adopting a “comprehend-first, transmit-later” approach, semantic communication significantly reduces transmitted data volume, minimizes bandwidth consumption, and improves communication efficiency. However, when semantic communication is tailored to specific tasks, receivers predominantly rely on task-relevant semantic information, rendering the transmission and reconstruction of complete semantic data redundant and inefficient.
Task-oriented semantic communication addresses this limitation by directly optimizing performance for specific tasks. It enables the transmitter to extract only task-relevant semantic information and allows the receiver to focus on enhancing task-specific outcomes [16]. Resource allocation in such systems is tightly linked to task goals, making them particularly advantageous for resource-constrained environments or tasks with strict real-time requirements [17]. Studies have demonstrated the efficiency and robustness of task-oriented semantic communication in various scenarios, including image retrieval [18] and federated learning [19], as well as considering resource-constrained devices [20] and MIMO channels [21].
Despite its advantages, task-oriented semantic communication introduces new security risks. Attackers may exploit transmitted semantic features or intermediate representations to infer the original input data through model inversion attacks, leading to potential privacy leakage [22]. While transmitting only task-relevant information offers some level of privacy protection, adversaries can still reconstruct raw data or extract sensitive information using model inversion techniques [23]. Traditional privacy protection methods face significant challenges in the context of semantic communication. For instance, differential privacy [24] adds perturbations to model inputs or outputs which degrade semantic communication performance [25], while federated learning protects privacy through parameter sharing, but it is vulnerable to gradient leakage attacks [26].
Adversarial training methods based on information bottleneck have been proposed to reduce the pixel-level quality of reconstructed images [27, 28, 29]. However, these methods often evaluate privacy protection using image quality metrics like peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM), which do not directly reflect the semantic fidelity of the reconstructed information. For example, images with low SSIM may still retain critical semantic features, while high PSNR does not necessarily correlate with improved task performance [30]. This underscores the need to unify evaluation metrics for both attackers and recipients, focusing on semantic fidelity [31, 16] rather than image quality.
In this paper, we introduce a diffusion-based semantic communication system with model inversion attack involved. Diffusion models, as a generative artificial intelligence technique, offer potential solutions to these challenges. They can generate high-quality data (e.g., images, audio) through a gradual denoising process, recovering the original data distribution from random noise via reverse diffusion [32, 33]. By incorporating Gaussian white noise in the forward process and gradually denoising during sampling, diffusion models have demonstrated adaptability to additive Gaussian noise channels. Recent studies have explored their integration with semantic communication for tasks like audio restoration [34] and image denoising [35, 36, 37], primarily relying on conditional diffusion models equipped with receiver-side guidance.
Diffusion models excel in recovering and reconstructing semantic information due to their progressive denoising mechanism and ability to mimic underlying data distributions. Moreover, they can enhance privacy protection by allowing the sender to regulate distortion levels, thereby constraining adversarial behavior. However, applying diffusion models to privacy-sensitive semantic communication introduces unique challenges. Conditional diffusion models often assume the availability of guidance, which may not always be feasible. For instance, the sender may lack the capability to provide guidance due to performance limitations, or the receiver may need to keep task-specific details confidential. Additionally, in classification tasks, the guidance itself often overlaps with the basis for decision-making, creating a conflict between providing guidance and completing the task. These constraints underline the need for innovative approaches to leverage diffusion models in privacy-aware semantic communication systems.
In this paper, we propose a framework, named DiffSem, that leverages diffusion processes to support privacy-sensitive task-oriented semantic communication. Inspired by the forward process of diffusion models, we employ an adaptive noise injection mechanism, named self-noising, to regulate the quality of signal from the transmitter. Diffusion-based U-Net is applied at the receiver to enhance task performance, with the option to incorporate self-referential label embedding for further improvements. Inspired by the concept of semantic fidelity at the receiver [16], we propose a novel evaluation metric to more accurately quantify the semantic information obtained by adversaries through model inversion attacks. By introducing a strong classifier trained to evaluate adversaries’ reconstructed images, our framework provides a direct assessment of the semantic fidelity achieved by adversaries while highlighting the inadequacy of traditional image quality metrics such as PSNR and SSIM in evaluating attack severity. Experimental results on the MNIST and CIFAR10 datasets demonstrate that DiffSem significantly improves receiver classification accuracy and achieves robust performance gains even under dynamic channel conditions.
Moreover, our findings reveal that traditional adversarial training often compromises the semantic fidelity of legitimate receivers without significantly reducing the semantic information extracted by adversaries, despite a marked deterioration in the quality of adversarial outputs. In contrast, our system implements a non-adversarial privacy protection strategy that enhances and preserves the semantic fidelity of legitimate receivers while restricting the fidelity attainable by adversaries. By conceptualizing adversaries as syntactic recovery agents and legitimate receivers as task-related semantic recovery agents, we explore the nuanced differences between syntactic and semantic information recovery. The self-denoising module in our system further helps refine this distinction by dynamically adjusting system outputs to optimize privacy protection and task performance. This unified framework not only advances task-oriented semantic communication but also provides a novel perspective on balancing task optimization and privacy protection in semantic communication systems.
In general, our contribution can be concluded as follows.
  • We implement a diffusion-based semantic communication framework, named DiffSem, which leverages the unique properties of diffusion processes to optimize semantic information reconstruction at the receiver, and incorporates self-referential label embedding to significantly improve task performance. This framework ensures high-quality semantic recovery while providing robust performance under dynamic channel conditions. Experimental results on the MNIST dataset demonstrate that DiffSem boosts the receiver’s classification accuracy by $10.03\%$ under high distortion conditions, while ensuring robust performance.
  • We develop a non-adversarial privacy protection strategy that focuses on maintaining and enhancing the semantic fidelity of legitimate receivers while strictly limiting the upper bound of semantic fidelity attainable by adversaries. Our approach balances privacy protection and task performance by carefully regulating the semantic information distribution of system outputs. Additionally, the system employs an adaptive noise injection mechanism to further distinguish between semantic and syntactic information recovery, offering a novel perspective on privacy protection in semantic communication.
  • We propose a new evaluation metric to accurately quantify the semantic information extracted by adversaries through model inversion attacks. By introducing a strong classifier to evaluate the semantic fidelity of reconstructed outputs, we provide a more precise and task-relevant framework for assessing privacy risks in semantic communication systems. In addition, we highlight a significant discrepancy between traditional image quality metrics (e.g., PSNR and SSIM) and task-relevant semantic information, underscoring the necessity of adopting semantic fidelity as a more appropriate evaluation criterion.
@@ II System architecture
In this section, we will first introduce the architecture of our system, then we will discuss the model inversion attack that can estimate the input data. Next, we will discuss the general evaluations of privacy preservation in our system model.
@@ II-A Diffusion based Task-oriented Semantic System
As shown in Fig. 1(a), assume there is a task-oriented transmitter-receiver pair that implements a specific task using the input raw data $x$ to get target object $y$. First, $x\in X$ is sent into a feature extractor module and a channel coding module to get neural-coded features $f\in F$. In our system, these two modules are both implemented by neural networks, and they can be regarded as a whole module named encoder $\Phi$ with parameter set $\phi$, like any other neural-network based joint-source-channel-coding (JSCC) system in related literature. Next, a self-noising module $\Psi$ is utilized to add a certain level of noise, based on system requirements, to get $z$ that is ready to transmit. These steps can be expressed as
@@
After transmitting over the air, the receiver $\mathcal{R}$ obtains
$\hat{z}=hz+\epsilon,$ (2)
where $h$ is channel gain, $\epsilon$ is channel noise that has standard variation $\sigma$ in the wireless channel $\mathcal{C}$. Then, $\mathcal{R}$ performs its task in two steps. First, it uses a diffusion module $\Delta$ to perform denoising process that generates the estimated feature $\hat{f}$. Then $\hat{f}$ is channel decoded and is used to perform specific task. These two functions are also combined into one module called classifier $\Theta$, just analogous to the encoder module introduced above. The process of $\mathcal{R}$ can be defined as
$\hat{y}=\mathcal{R}_{(\delta,\theta)}(\hat{z})=\Theta_{\theta}(\Delta_{\delta}(\hat{z}))=\Theta_{\theta}(\hat{f}),$ (3)
where $\delta$ and $\theta$ are the parameter set in denoising module $\Delta$ and classifier module $\Theta$, respectively, and $\hat{y}$ is the task output. Note that in wireless transmissions, $z,\hat{z}\in\mathbb{C}^{\kappa}$, where $\kappa$ means complex channel use. Since neural network only process real numbers, without further notice, let $k=2\kappa$, and the length of $z$ ($f$ as well) is $k$ in our system’s implementation.
There are two premises that need to be explained here. First, we assume that there is a trustworthy third party that can define the specific task and arrange models in each node of the system. Thus, we define that $\Theta$ is reachable only in $\mathcal{R}$ when models are trained and deployed by the third-party, which is a common prerequisite in task-oriented communications. Moreover, we argue that neither the plaintext nor the encrypted forms of $x$ and $y$ should be transmitted directly. Specifically, due to the limited computational capability of the transmitter, even computing $y$ locally is computationally challenging. As for $x$, we assume that both the receiver and potential eavesdroppers are curious and may attempt to infer information about $x$. Consequently, transmitting $x$—even in encrypted form—may still enable the receiver to recover $x$, unless extremely inefficient privacy-preserving inference mechanisms are employed. Therefore, to ensure data privacy, direct transmission of any form of $x$ or $y$ should be strictly avoided.
@@ II-B Model inversion attack
Model inversion attack is an adversarial technique that targets machine learning models. The adversary can query the model and analyze how the model responds to attempt to infer the original input. It poses significant risks to data privacy, especially when sensitive information is involved.
As shown in Fig. 1(b), an adversary $\mathcal{A}$ may want to perform model inversion attack on the transmitter by recovering the estimation of $x$ (denoted by $\hat{x}$), given the detected $\hat{\dot{z}}$ from a different wireless channel path. $\mathcal{A}$ tries to recover the estimation of $x$, denoted by
$\hat{x}=\mathcal{A}_{a}(\hat{\dot{z}})=\mathcal{A}_{a}(\dot{h}z+\dot{\epsilon}),$ (4)
where $\dot{h}$ and $\dot{\epsilon}$ are channel gain and noise in the adversary’s channel, respectively, and $a$ is the parameter set of $\mathcal{A}$.
Assume $\mathcal{A}$ can access $\mathcal{T}$ repeatedly on the target device but cannot obtain the parameters and training data, and $\Theta$ is confidential to $\mathcal{A}$ and will not be accessed in any cases. Generally, the adversary first prepares a dataset that has a similar distribution as $X$, named $\dot{X}$, and sends $\dot{x}$ into $\mathcal{T}$ to get $\hat{\dot{z}}$. When it gets enough data pairs $\{\dot{x},\hat{\dot{z}}\}$, it trains on the data pairs using mean squared error (MSE) loss
$\mathcal{L}_{\text{MSE}}(x_{i},\hat{x}_{i})=\frac{1}{M}\sum^{M}_{i}\|x_{i}-\hat{x}_{i}\|^{2},$ (5)
to minimize the reconstruction distortion. This is called a black-box setting [38].

### v2

# Diffusion-aided Task-oriented Semantic Communications with Model Inversion Attack

@@ Abstract
Semantic communication enhances transmission efficiency by conveying semantic information rather than raw input symbol sequences. Task-oriented semantic communication is a variant that tries to retains only task-specific information, thus achieving greater bandwidth savings. However, these neural-based communication systems are vulnerable to model inversion attacks, where adversaries try to infer sensitive input information from eavesdropped transmitted data. The key challenge, therefore, lies in preserving privacy while ensuring transmission correctness and robustness. While prior studies typically assume that adversaries aim to fully reconstruct the raw input in task-oriented settings, there exist scenarios where pixel-level metrics such as PSNR or SSIM are low, yet the adversary’s outputs still suffice to accomplish the downstream task, indicating leakage of sensitive information. We therefore adopt the attacker’s task accuracy as a more appropriate metric for evaluating attack effectiveness. To optimize the gap between the legitimate receiver’s accuracy and the adversary’s accuracy, we propose DiffSem, a diffusion-aided framework for task-oriented semantic communication. DiffSem integrates a transmitter-side self-noising mechanism that adaptively regulates semantic content while compensating for channel noise, and a receiver-side diffusion U-Net that enhances task performance and can be optionally strengthened by self-referential label embeddings. Our experiments demonstrate that DiffSem enables the legitimate receiver to achieve higher accuracy, thereby validating the superior performance of the proposed framework.
@@ I Introduction
Deep learning based semantic communication systems have emerged as a promising solution to enhance information transmission efficiency in sixth-generation (6G) wireless networks. Unlike traditional paradigms that prioritize the accurate symbol-level transmission, semantic communication shifts the focus to conveying the intended meaning through symbols [1]. In such systems, the transmitter and receiver collaboratively design mappings between meanings and channel symbols, naturally aligning with the principles of joint source-channel coding (JSCC) [2, 3]. This approach accommodates more complex data sources (e.g., images) and diverse distortion measures, such as classification loss [4] or task-oriented metrics [5]. As a result, semantic communication has been studied for various types of sources such as text [6, 7], visual data [8, 9, 10], and video [11, 12], and has also been applied in edge computing [13, 14] and autonomous driving [15, 16]. By adopting a “comprehend-first, transmit-later” approach, semantic communication significantly reduces transmitted data volume, minimizes bandwidth consumption, and improves communication efficiency.
As a variant of semantic communication, task-oriented semantic communication shifts the target from reconstructing the transmitter’s raw input to directly enabling the receiver to accomplish a given task, thereby allowing the transmitter to further compress the transmitted content [17, 18]. For example, in a typical task-oriented semantic communication scenario targeting the image classification task, the sender sends the semantic representation of the picture, and the receiver uses the semantic information to perform the classification task directly, skipping the process of recovering the picture. Studies have demonstrated the efficiency and robustness of task-oriented semantic communication in various scenarios, including image retrieval [19] and federated learning [20], as well as scenarios involving resource-constrained devices [21] and MIMO channels [22].
Despite its advantages, task-oriented semantic communication, like all neural-based systems, is vulnerable to model inversion attacks, in which an adversary eavesdrops on the transmitter’s output and attempts to reconstruct sensitive information of the input [23]. Conventional privacy-preserving techniques, however, are difficult to apply in this context. For example, differential privacy perturbs model inputs or outputs [24], but such perturbations can significantly degrade semantic communication performance [25]. Similarly, federated learning avoids direct data sharing by transmitting parameters, yet remains vulnerable to gradient leakage attacks [26].
Some existing works use adversarial training methods, inspired by the information bottleneck theory, that aim to reduce the pixel-level quality of reconstructions [27, 28, 29], using image-quality metrics such as peak signal-to-noise ratio (PSNR) or structural similarity index measure (SSIM). From an information-theoretic perspective, requiring an adversary to fully reconstruct the original data is not well-founded, as the transmitter never conveys sufficient information to enable exact recovery. In fact, preventing input reconstruction merely requires constraining the transmitted information to contain only task-relevant features, without the need for additional countermeasures.
Motivated by this observation, we consider a generalized notion of model inversion attacks: an attack is deemed successful if the adversary can reconstruct sensitive information that is sufficient to accomplish the task. For classification tasks, this definition corresponds to evaluating the accuracy of the adversary’s reconstruction under a strong classifier [17], similar to the concept of “semantic fidelity” in [17, 30]. We revisit part of the aforementioned work and argue that their proposed defenses [29, 31] are in fact weak. Even when pixel-level reconstructions appear poor, an adversary may still succeed if the recovered semantics are sufficient for the task. Conversely, restricting task-agnostic recoverability while preserving task-specific semantics provides a principled means of protecting privacy, which is aligned with the analysis in [32]. Building on this perspective, we revisit the notion of privacy: in task-oriented systems, we argue that privacy should be measured by the performance gap between the legitimate receiver and the adversary. In other words, the source of privacy lies in the information asymmetry between them.
This paper proposes DiffSem, a diffusion-aided task-oriented semantic communication system which enhances task performance while limiting adversarial recovery. A diffusion model is a generative architecture that learns to recover the original data distribution from noise through forward noising and reverse denoising [33, 34]. Such models have been applied to tasks such as audio restoration and image denoising within semantic communication, where conditional diffusion and receiver-side guidance are typically employed to help the receiver refine its reconstructions [35, 36, 37, 38]. However, in privacy-preserving semantic communication systems, the receiver’s task can be confidential, that is, neither the transmitter nor potential adversaries are allowed to access task information [39]. To make sure that DiffSem works in the confidential scenario, it couples a transmitter-side self-noising module that adaptively regulates how much task-specific feature is observable while compensating for channel noise, with a guidance-free receiver diffusion U-Net that enhances task performance under dynamic channels; the receiver can be further strengthened by self-referential label embeddings while remaining guidance-free. Conceptually, the self-noising module exploits the semantic-information mismatch: it reduces task-agnostic recoverability from the transmitted features without depriving a task-aware receiver of the semantics needed for the downstream task.
In general, our contributions can be summarized as follows.
  • We generalize the notion of model inversion attack by defining its success in terms of the attacker’s task accuracy rather than pixel-level similarity. Experiments show that even when reconstructions have low PSNR or SSIM, attackers can still achieve high task accuracy, indicating that conventional defenses focused on image pixel-level quality are largely weak.
  • We develop a non-adversarial privacy mechanism that exploits the intrinsic mismatch between the semantics needed by the legitimate task and the broader content an attacker must reconstruct. By limiting task-agnostic recoverability while preserving task-specific semantics, our approach mitigates model inversion without the utility loss associated with adversarial training.
  • We propose DiffSem, a diffusion-aided semantic communication framework for scenarios where the receiver’s task is confidential. DiffSem introduces a transmitter-side self-noising mechanism that adaptively regulates the radiated semantic information to compensate for channel noise, and a guidance-free receiver-side diffusion U-Net that enhances task performance. An optional self-referential label embeddings further improves performance without violating the guidance-free constraint. Our experiments on MNIST and CIFAR-10 datasets show that DiffSem significantly increases the legitimate receiver’s accuracy.
@@ II System architecture
In this section, we first describe the overall system architecture and then analyze model inversion attacks as a representative privacy threat. Finally, we introduce the evaluation framework used to quantify the privacy guarantees of our design.
@@ II-A Diffusion based Task-oriented Semantic System
As shown in Fig. 1(a), assume there is a task-oriented transmitter-receiver pair that performs a specific task using the input raw data $x$ to get target output $y$. First, $x\in X$ is sent into a feature extractor module and a channel coding module to get neural-coded features $f\in F$. In our system, these two modules are both implemented by neural networks, and they can be regarded as a whole module named encoder $\Phi_{\phi}$ with parameter set $\phi$, like any other neural-network based joint-source-channel-coding (JSCC) system in related literature. Next, a self-noising module $\Psi$ is utilized to add a certain level of noise, based on system requirements, to get $z$ that is ready to transmit. These steps can be expressed as
@@
Assuming additive white Gaussian noise (AWGN) channel is used [^1] Our model also covers block fading channel[40] with receiver’s channel state information (CSI). Consider $x$ being transmitted within a block, a one-tap coherent equalizer maps $\hat{z}=hz+\epsilon$ to $z+\tilde{\epsilon}$ where the noise variance of $\tilde{\epsilon}$ is rescaled by $1/|h|^{2}$ compared to $\epsilon$ [8, 41, 42]. We leave a detailed treatment of fading channels to our future work.. After transmitting over the air, the receiver first observes the complex-baseband signal
$\hat{z}=z+\epsilon,$ (2)
where $\epsilon\!\sim\!\mathcal{N}(0,\sigma^{2}\mathbf{I})$ is additive white Gaussian noise (AWGN) in the wireless channel $\mathcal{C}$. Then, $\mathcal{R}$ performs its task in two steps. First, it uses a diffusion module $\Delta_{\delta}$ to get the estimated feature $\hat{f}$ through denoising; second, $\hat{f}$ is channel decoded and is used to perform a specific task. Decoding and task execution are also combined into a single module, referred to as the classifier $\Theta_{\theta}$, analogous to the encoder module introduced above. Hence the receiver is defined as
[…]
where $\delta$ and $\theta$ are the parameter sets in denoising module $\Delta_{\delta}$ and classifier module $\Theta_{\theta}$, respectively, and $\hat{y}$ is the task output. Note that in wireless transmissions, $z,\hat{z}\in\mathbb{C}^{\kappa}$, where $\kappa$ means complex channel use. Since neural networks only process real numbers, without further notice, let $k=2\kappa$, and the length of $z$ ($f$ as well) is $k$ in our system’s implementation.
Remarks. We assume that there is a trustworthy third party that can define the specific task and arrange models in each node of the system. Thus, we define that $\Theta_{\theta}$ is reachable only in $\mathcal{R}$ when models are trained and deployed by the third-party, which is a common prerequisite in task-oriented communications [43].
@@ II-B Model Inversion Attack
Model inversion attack, proposed by Fredrikson et al. [44], is a representative threat model in which an adversary attempts to reconstruct sensitive input information from the outputs of a machine learning model. In semantic communication systems, the concern about model inversion attack has been explicitly raised in [29]. In their setting, an adversary continuously queries the semantic transmitter to obtain noisy features, and then trains an inversion model to recover private inputs. They adopt a black-box adversarial assumption and employ mean squared error (MSE) as the reconstruction metric. Motivated by these findings, we now formalize the threat model for JSCC-based semantic communications under the following assumptions [45]:
  • Black-box access. The adversary $\mathcal{A}_{a}$ has no access to the model parameters or the training dataset of the transmitter $\mathcal{T}$.
  • Distribution mismatch. Unlike the legitimate transmitter, $\mathcal{A}_{a}$ does not possess data drawn from the same distribution as $X$, and must construct a surrogate dataset $\dot{X}$ from a related but mismatched distribution.
  • Chosen-plaintext attack. $\mathcal{A}_{a}$ may freely construct input samples $\dot{x}\in\dot{X}$, obtain the corresponding post-channel outputs $\hat{\dot{z}}$ delivered by $\mathcal{T}$, and collect sufficient pairs $\{\dot{x},\hat{\dot{z}}\}$ to train itself, as shown in Fig. 2.
We also employ the wireless settings similar to [12], such that $\mathcal{A}_{a}$ can only eavesdrop on the communication link rather than directly observe the output of the model. As shown in Fig. 1(b), $\mathcal{A}_{a}$ observes $\hat{\bar{z}}$ on a separate channel and attempts to reconstruct the transmitted symbol $x$. This is modeled as
$\hat{x}=\mathcal{A}_{a}(\hat{\bar{z}})=\mathcal{A}_{a}\big(\bar{z}+\dot{\epsilon}\big),$ (4)
where $\dot{\epsilon}\!\sim\!\mathcal{N}(0,\dot{\sigma}^{2}\mathbf{I})$ denotes the noise in the adversary channel, and $a$ is the parameter set of $\mathcal{A}_{a}$. To train its inversion model, $\mathcal{A}_{a}$ minimizes the reconstruction distortion using the MSE loss:
$\mathcal{L}_{\text{MSE}}(x_{i},\hat{x}_{i})=\frac{1}{M}\sum_{i=1}^{M}\|x_{i}-\hat{x}_{i}\|^{2}.$ (5)